# Hyperparameters are all you need: Using five-step inference for an original diffusion model to generate images comparable to the latest distillation model

### v2

@@ Abstract
The diffusion model is a state-of-the-art generative model that samples images by applying a neural network iteratively. However, the original sampling algorithm requires substantial computation cost, and reducing the sampling step is a prevailing research area. To cope with this problem, one mainstream approach is to treat the sampling process as an algorithm that solves an ordinary differential equation (ODE). Our study proposes a training-free inference plugin compatible with most few-step ODE solvers. To the best of my knowledge, our algorithm is the first training-free algorithm to sample a 1024 x 1024-resolution image in 6 steps and a 512 x 512-resolution image in 5 steps, with an FID result that outperforms the SOTA distillation models and the 20-step DPM++ 2m solver, respectively. Based on analyses of the latent diffusion model’s structure, the diffusion ODE, and the Free-U mechanism, we explain why specific hyperparameter couplings improve stability and inference speed without retraining. Meanwhile, experimental results also reveal a new design space of the latent diffusion ODE solver. Additionally, we also analyze the difference between the original diffusion model and the diffusion distillation model via an information-theoretic study, which shows the reason why the few-step ODE solver designed for the diffusion model can outperform the training-based diffusion distillation algorithm in few-step inference. The tentative results of the experiment prove the mathematical analysis.
@@ 1 Introduction
The diffusion probability model (DM) proposed by Ho et al. (2020a), a state-of-the-art image generation model with solid mathematics instruction on image sampling, is among the most popular research areas in computer vision. A typical diffusion model, such as stable diffusion (SD) (Rombach et al. (2022)), is inspired by the Brownian diffusion process in physics. It consists of two different processes: diffusion and reverse. The diffusion process adds small variants of Gaussian noise to the image step by step to convert an image into a pure Gaussian noise, while the reverse process uses the score matching algorithm (Song et al. (2020b)) to train a neural network studying the gradient of the diffusion process by denoising noisy images (Luo (2022)).
Like the Brownian process in physics, the diffusion process can be described by a diffusion stochastic differential equation (SDE). Moreover, the research proposed by Song et al. (2020b) shows that there exists an ordinary differential equation (ODE) to describe the reverse process. And using the existing mathematical tool for ODE solving enhances the capability of the diffusion model to sample an image.
However, compared to typical algorithms such as generative adversarial networks and variational autoencoders, solving diffusion ODEs still requires substantial computational resources. Hence, coping with this problem remains an active research area.
@@
Research related to it can be categorized into two classes. The first class proposes a training-free algorithm aimed at providing a good SDE and ODE solver tailored to the diffusion process (Ho et al. (2020a); Lu et al. (2025); Song et al. (2020a); Zhao et al. (2023); Karras et al. (2022); Liu et al. (2022)). The second class deems the whole inference process, which uses a neural network iteratively, as a large neural network, and the algorithm reducing the inference steps is equivalent to train a small distillation model (Luo et al. ; Song & Dhariwal (2023); Lin et al. (2024); Yin et al. (2024b; a); Salimans & Ho (2022); Sauer et al. (2024); Kim et al. (2023); Song et al. (2023); Liu et al. (2023)). The training-free method typically necessitates a 20-step inference, like the DPM++ solver proposed by Lu et al. (2025), while algorithms requiring additional training can generate an image within four to eight steps. However, our results show that by selecting the appropriate hyperparameter for the ODE solver and adding a training-free diffusion model decorator to leverage the latent diffusion model’s capabilities, the inference speed can exceed expectations.
@@
To that end, we propose a new inference method that generates 512 x 512 images in 5 steps without additional training, and its Frechet Inception Distance (FID) (Heusel et al. (2017)) performance surpasses both the DPM++ 2m solver and the SOTA distillation model in 20 steps and 8 steps, respectively. We also apply our algorithm to 1024 x 1024 image generation, and the FID performance of the eight-step and six-step inference exceeds most of the latest distillation models. We perform extensive experiments using the COCO 2014 (Lin et al. (2014)), COCO 2017 (Lin et al. (2014)), and LAION datasets (Schuhmann et al. (2022)), and exhibit the generation result in Fig 1. We also use information theory to explain why our algorithm achieves a strong FID result.
@@
Figure 1: The comparison between our algorithm and UniPC solver in few step inference
@@ 2 Related Work
2.1 Diffusion ODE
The image-generation process of the diffusion model can be deemed as solving a special diffusion ODE. A typical diffusion model has two processes: forward and reverse. The forward process perturbs the image by adding Gaussian noise to it, which can be described by the following SDE:
$dx=f(t)xdt+g(t)dw,$ (1)
where the $w$ is a standard Wiener process. Choosing different $g(t)$ and $f(t)$ corresponds to applying different algorithms. In this study, we use the variant preserve (VP) SDE proposed by Song et al. (2020b) to describe the diffusion process, which corresponds to the denoised diffusion probability model (DDPM) algorithm (Ho et al. (2020b)). Specifically, its choice is as follows:
$f(t)=-\frac{1}{2}\beta(t)=\frac{d\log{\alpha_{t}}}{dt},$ (2)
$g(t)=\sqrt{\beta(t)},$ (3)
$\beta(t)=\beta_{min}+t(\beta_{max}-\beta_{min}).$ (4)
Explicitly, in the diffusion process mentioned above, the variant of noise added at step $t$ is $\beta_{t}$. Explicilty, $\beta_{max}$ and $\beta_{min}$ can be determined arbitrarily if it can ensure $p(x_{T}|x_{0})$ is a standard Gaussian distribution. In the timestep $t$, the noisy feature $x_{t}$ with given $x_{0}$ belongs to a Gaussian distribution described in the eq.˜5.
$p(x_{t}|x_{0})=\mathcal{N}(\alpha(t)x_{0};\sigma^{2}(t)I).$ (5)
Typically, to transform an image into a standard Gaussian noise, the DDPM algorithm requires $T=1000$ steps, and $\alpha_{T}$ is closed to the zero. Additionally, using the reparameterization algorithm, we can sample $x_{t}$ in a single step (Ho et al. (2020a)).
In contrast, the reverse process, which corresponds to generating an image from a pure Gaussian noise, can be described by the following time-dependent ODE proposed by Karras et al. (2022):
$\frac{dx}{dt}=\frac{\dot{s}(t)}{s(t)}x_{t}+s(t)^{2}\dot{\sigma}(t)\epsilon_{\theta}(x_{t},t,c),$ (6)
where $c$ is the condition of input, like the caption that describes the image, and the $\epsilon_{\theta}$ is modeled by a neural network. The $\epsilon_{\theta}$ has a relation between the score function, which is described in the eq.˜7:
$\epsilon_{\theta}(x_{t},t,c)=-\sigma(t)score_{\theta}(x_{t},c),$
$score_{\theta}(x_{t},c)=\nabla_{x_{t}}\log{p_{\theta}(x_{t},c)},$ (7)
And the $p_{\theta}(x_{t},c)$ is the probability distribution of noisy images $x_{t}$ modeled implicitly by the neural network with parameters $\theta$. Moreover, the score function normally used in VP SDE is trained by eq.˜8
[…]
and $q_{t}(x)$ is the probability distribution of the random variables sampled by adding noise to the training images in timestep $t$, and $q_{0}$ means the distribution of clean images.
In the eq.˜6, the $s(t)$ and $\sigma(t)$ is defined by eq.˜9 and eq.˜10:
$\sigma(t)=\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}-1},$ (9)
$s(t)=1/\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}},$ (10)
where $\beta_{d}=\beta_{max}-\beta_{min}$.
Deeming the loss of the DDPM algorithm as a score function described by the eq.˜8 can help exploit the advantage of adapting the original diffusion model to the common diffusion ODE framework proposed by Karras et al. (2022).
2.2 Diffusion ODE Solver
Sampling an image from a diffusion model is equivalent to solving the eq.˜6. Via the research from Lu et al. (2022), when given a noisy input $x_{s}$ and a target timestep $t$, the correct noisy feature output $x_{t}$ can be gained by solving the eq.˜11 :
@@ 2.2 Diffusion ODE Solver
where $\lambda_{t}$ is the log Signal-to-Noise Ratio (SNR) of the noisy feature $x_{t}$.
Sometimes, to cope with the high guidance scale on the classifier free guidance algorithm (CFG) (Ho & Salimans (2022)), researchers tend to replace eq.˜11 with the eq.˜12 to get a better result (Lu et al. (2025)):
$x_{t}=\frac{\sigma({t})}{\sigma(s)}x_{s}-\sigma_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{\lambda}\hat{x}_{\theta}(x_{t(\lambda)},\lambda)d\lambda,$ (12)
where $\hat{x}_{\theta}$ is a clean feature predicted indirectly by the neural network.
By using the Taylor Expand in the $\epsilon_{\theta}$ or $x_{\theta}$ (Lu et al. (2022; 2025)), we get the following eq.˜13:
@@
However, calculating the neural networks’ deviation requires substantial computational resources. Researchers typically use the Runge-Kutta method (Lu et al. (2025); Zhao et al. (2023)) to approximate the deviation.
Explicitly, the DPM++ 2s algorithm approximate $x_{t}$ via the following formula:
$x_{t}$ $=\frac{\sigma({t})}{\sigma(s)}x_{s}-\alpha_{t}(e^{-h_{t}}-1)\hat{x}_{\theta}(x_{s},s)$
$+\frac{1}{2r_{t}}*\alpha_{t}(e^{-h_{t}}-1)*(\hat{x}_{\theta}(\mu,m)-\hat{x}_{\theta}(x_{s},s)),$ (14)
in which $r_{t}$ can be selected arbitarily between 0 and 1, and $h_{i}$, $m$ and $\mu_{i}$ is given below:
$h_{t}$ $=\lambda_{t}-\lambda_{s},$ (15)
$r_{t}$ $=\frac{\lambda_{m}-\lambda_{s}}{h_{i}},$ (16)
$\mu_{t}$ $=\frac{\sigma_{m}}{\sigma_{s}}x_{s}-\alpha_{m}(e^{-r_{t}h_{t}}-1)\hat{x}_{\theta}(x_{s},s),$ (17)
To get the first-order deviation of the $\epsilon_{\theta}$, the number of function evaluations (NFE) in DPM++ 2s is 2. While DPM++ 2m utilizes the latest cached output in the previous inference to calculate the approximation of the first-order deviation, and hence only requires 1 NFE, which is a typical multistep algorithm in ODE solving. Simultaneously, the UniPC and DPM V3 solver cached all of the previous output to correct and predict a high-order deviation.
The eq.˜6 describes a continuous curved trajectory, and the ODE solver approximates it by discretizing the path with a series of lines. Research on reverse ODE solvers aims to precisely estimate the next step, thereby minimizing truncation error from pool discretization in the few-step inference. On the other hand, the discrete method proposed by Karras et al. (2022) does not aim at design a new ODE solving algorithm, but proposes a choice of moving distance $h_{t}$ to enhance the sample quality. Via the analysis of the Karras’ research, the $h_{t}$ should be large at the beginning of the inference, and then decrease monotonically.
@@ 2.3 Diffusion Distillation
The training-free method optimizes the truncation error at each step and thus minimizes the need for additional inference steps for the precise discretization. Unlike this method, the distillation algorithm (Lin et al. (2024); Sauer et al. (2024); Luo et al. ; Song & Dhariwal (2023); Salimans & Ho (2022)) considers the whole inference process of applying a neural network iteratively as an extensive neural network. The distillation algorithm, which typically trains a small neural network, reduces its inference steps. However, training a neural network with high-quality content, fast inference, and no additional parameters is difficult. Thus, the distillation algorithm, through the analysis of information theory (Li & Zhang (2025)) and experiments, can sometimes reduce the diversity of the generated images, and affect the FID performance. This approach may pose a potential risk in the future. AI training requires the use of Internet data, and in some countries, like Japan, all the data can be used in AI training by obeying a series of basic rules. However, when a human draws an image via their craft, they will sometimes be affected by the image they have previously seen. If they generate a similar image unconsciously, without directly using a similar work as a reference during creation, such a generation will not be deemed plagiarism. Compared to this, generative models trained on Internet images are often criticized, with people claiming that if those models generate content similar to their dataset, it should be considered plagiarism. In such a situation, diversity is a metrics that affect people’s impression of generative content.
@@ 2.4 Latent Diffusion Model
Another method for reducing the computational requirements of inference is to encode images into a low-dimensional latent space. Stable diffusion performs inference in a 64 x 64 x 4 latent space to generate a 512 x 512 x 4 PNG image. One experiment from Rombach et al. (2022) shows that the semantics of the image, which rely on the long-range relationship between different parts of the images, can be modeled by an attention mechanism (Vaswani et al. (2017)). In contrast, its perceptual information, which provides the image with sufficient real details, is strongly tied to local regions and can be modeled by a convolutional neural network (CNN) (e.g. Li et al. (2021)). Hence, the stable diffusion model first encodes the image into a low-dimensional latent space via a $\beta$-variational autoencoder (Burgess et al. (2018)) to let the following inference ignore the perceptual detail. Then, it performs diffusion inference in latent space via a U-Net (Ronneberger et al. (2015)) with an attention mechanism to model the semantic information. Experiments show that compared with vector quantify VAE (Wu & Flierl (2020)), the normal $\beta$-VAE with a small KL-term constraint generates higher-quality images.
@@ 3 Approach
3.1 Algorithm Based on Truncation Error Analysis
To design a training-free, few-step inference algorithm, we should first consider the meaning of the original diffusion models’ output, the score function. And then compare it with the distillation model’s output.
By rewriting the eq.˜8 as follows:
$\mathcal{L}_{\theta}=\frac{E_{x_{0}}[q(x_{t}|x_{0})||s_{\theta}-\nabla_{x}\log q(x_{t}|x_{0})||^{2}]}{E_{x_{0}}[q(x_{t}|x_{0})]},$ (18)
We find out that the score function is pointing to the average of all possible images that can cause this noisy image. In an extreme few-step inference case, like one or two NFE inference, the final output will be blurred, since in that step, the score function only points to the centroid of images. Meanwhile, the distillation algorithm, like rectify flow proposed by Liu (2022), will tend to modify the output and let it point to a special sample.
Although the task of modifying the score function to allow it to point to a special sample without additional training is hard, one can presume that the score function $\nabla_{x}\log q(x_{t})$ can be approximated by the conditional score function $\nabla_{x}\log q(x_{t}|x_{0})$ in the least noisy stage, since the possible images are very limited. In this situation, the inference task can be approximated by a pure denoising task, and we do not need further inference if we can modify the U-Net to perform the denoising task without additional training.
Since the changing $\nabla_{x}\log p_{\theta}(x_{t})$ to the $\nabla_{x}\log p_{\theta}(x_{t}|x_{0})$ means removing the blurred part of the result caused by the average, one of the natural method is to enhance the denoise capability of U-Net, which can be achieved by modifying the skip connection and the backbone feature of the U-Net.
Previous study (Si et al. (2024)) adds a scalar, fast Fourier transformation (FFT) and inverse fast Fourier transformation (iFFT) modules to the skip connection, which the backbone feature being modified by the scalar can use to enhance denoising capability. Moreover, the skip connection being modified by the FFT and iFFT can prevent further smoothing and add details by amplifying the low-frequency part of the noisy backbone feature map. The decorator only applies to the first two skip connections. The scaler that affects the backbone feature is $b_{1}$ and $b_{2}$, while the scaler that affects the FFT and iFFT is $s_{1}$ and $s_{2}$.
However, we don’t know which step could be changed simply from the score function to the conditional score function, and one of the vanilla ways is to apply free-U at all time steps. This method has been seen as helpful to slightly modify the score function and make it more similar to the distillation’s output. The previous research from Ma et al. (2024) also shows that compared to the original diffusion model, the diffusion distillation model has small skip connection scalers, since the skip connection itself will abate the result in the block below that layer, while the few-step inference tends to excute a huge transportation from a standard Gaussian distribution to a complex probability distribution, and hence require the skip connection reduce its’ effect that abate the block below it. However, via the mathematical analysis of the diffusion ODE and the experimental result. This vanilla method in the extreme few-step inference of high-resolution images, such as 6 NFE, would not work properly.

### v3

# F-scheduler: illuminating the free-lunch design space for fast sampling of diffusion models

@@ Abstract
Diffusion models are the state-of-the-art generative models for high-resolution images, but sampling from pretrained models is computationally expensive, motivating interest in fast sampling. Although Free-U Net is a training-free enhancement for improving image quality, we find it ineffective under few-step (<10<10) sampling. We analyze the discrete diffusion ODE and propose F-scheduler, a scheduler designed for ODE solvers with Free-U Net. Our proposed scheduler consists of a special time schedule that does not fully denoise the feature to enable the use of the KL-term in the $\beta$-VAE decoder, and the schedule of a proper inference stage for modifying the U-Net skip-connection via Free-U Net. Via information theory, we provide insights into how the better scheduled ODE solvers for the diffusion model can outperform the training-based diffusion distillation model. The newly proposed scheduler is compatible with most of the few-step ODE solvers and can sample a 1024 x 1024-resolution image in 6 steps and a 512 x 512-resolution image in 5 steps when it applies to DPM++ 2m and UniPC, with an FID result that outperforms the SOTA distillation models and the 20-step DPM++ 2m solver, respectively.
Machine Learning, ICML
@@ 1 Introduction
The diffusion probability model (DPM) proposed by Ho et al. (2020) is a state-of-the-art image generation model. However, compared to typical algorithms (Goodfellow et al., 2020; Kingma & Welling, 2013), sampling from DPMs requires multiple neural network evaluations and incurs substantial computational cost. Hence, accelerating the sampling from DPMs remains an active research area.
@@
Research related to it can be categorized into two classes. The first class proposes a training-free algorithm aimed at providing a good SDE and ODE solver tailored to the diffusion process (Ho et al. (2020); Lu et al. (2025); Song et al. (2020); Zhao et al. (2023); Karras et al. (2022); Liu et al. (2022)). The second class deems the whole inference process as a large neural network, and reduces the inference steps by training a small distillation model (Luo et al. ; Song & Dhariwal (2023); Lin et al. (2024); Yin et al. (2024b, a); Salimans & Ho (2022); Sauer et al. (2024); Kim et al. (2023); Song et al. (2023); Liu et al. (2023)), which can generate images in four to eight steps.
Most training-free ODE solvers are not tailored to diffusion models, except the method of Karras et al. (2022), which shows that a uniform discretizing schedule causes increasing truncation error in few-step inference and proposes a non-uniform schedule with larger steps in high-noise regions and smaller steps in low-noise regions. However, this method is not customized for the latent diffusion model.
The latent diffusion model itself includes $\beta$-VAE and skip connections, both of which can be additionally boosted by a better scheduler in the design space. The $\beta$-VAE trained with a little noise can also reduce the inference budget by preventing the total denoising of images. However, adding same noise to the latent feature at different resolutions also has distinct effects(Hoogeboom et al., 2023), which necessitates the investigation of the design space. Furthermore, we find that the U-Net decorator, which aims to change the skip connection to enhance the denoising process, has the side effect of mimicking the distillation model(Ma et al., 2024) and changing the discretization result. We use ODE theory to fix the unwilling discretization effect, and use information theory to explain why the distillation model has the diversity problem, which provides an explanation of the phenomenon proposed by Bertrand et al. (2025), and proves the advantage of exploiting the design space of using a skip-connection decorator.
To that end, by exploring the design space for the latent diffusion model, we implement a new inference method compatible with most ODE solvers. Applied to UniPC(Zhao et al., 2023), it can generate 512 x 512 images in 5 steps without additional training, and its Frechet Inception Distance (FID) (Heusel et al. (2017)) performance surpasses both the DPM++ 2m solver and the SOTA distillation model in 20 steps and 8 steps, respectively, see Fig 1.
Figure 1: The comparison of FID performance using Free-U in different steps via ODE solvers in COCO 2014 with guidance scale $5.$ The size of the circle is proportional tothe amount of training cost.
@@
We also apply our algorithm to 1024 x 1024 image generation, and the FID performance of the eight-step and six-step inference exceeds most of the latest distillation models. We perform extensive experiments using the COCO 2014 (Lin et al. (2014)), COCO 2017 (Lin et al. (2014)), and LAION datasets (Schuhmann et al. (2022)), and exhibit the generation results in Fig 2.
@@
Figure 2: The comparison between our algorithm and UniPC solver in few step inference with the prompt ”an astronaut riding a horse on Mars”.
@@ 2 Related Work
2.1 Diffusion ODE and ODE Solver
The image-generation process of the diffusion model can be deemed as solving a special diffusion ODE. A typical diffusion model has two processes: forward and reverse. The forward process perturbs the image by adding Gaussian noise to it. Hence, in the timestep $t$, the probability density of the noisy feature $x_{t}$ with given $x_{0}$ belongs to a Gaussian distribution described in the Equation 1.
$q(x_{t}|x_{0})=\mathcal{N}(\alpha_{t}x_{0};\sigma^{2}_{t}I).$ (1)
where $\alpha_{t}$ and $\sigma_{t}$ can be choose arbitrarily if it can ensure the final $p(x_{T}|x_{0})$ close to a standard Gaussian distribution.
In contrast, the reverse process, which corresponds to generating an image from pure Gaussian noise, can be described by a time-dependent ODE. Via the research from Lu et al. (2022), when given a noisy input $x_{s}$ and a target timestep $t$, the correct noisy feature output $x_{t}$ can be gained by solving the Equation 2 :
@@ 2.1 Diffusion ODE and ODE Solver
where $\lambda_{t}$ is the log Signal-to-Noise Ratio (SNR) of the noisy feature $x_{t}$,and the $\epsilon_{\theta}$ is modeled by a neural network with ground truth proportional to the score function $\nabla_{x}\log{q(x_{t})}$, which is trained with the objective function described in Equation 3
$\mathcal{L}_{\theta}=E_{x_{0}}[q(x_{t}|x_{0})||\frac{\epsilon_{\theta}(x_{t},t)}{\sigma_{t}}-\nabla_{x}\log q(x_{t}|x_{0})||^{2}],$ (3)
By using the Taylor Expand in the $\epsilon_{\theta}$ (Lu et al. (2022, 2025)), we get the following Equation 4:
@@
where $h_{i}\propto\lambda_{t}-\lambda_{s}$
However, calculating the neural networks’ deviation requires substantial computational resources. Researchers typically use different algorithm, like the Runge-Kutta method (Lu et al. (2025); Zhao et al. (2023)), to approximate the deviation.
2.2 Time Schedule
The original diffusion ODE describes a continuous curved trajectory, and the Equation 2 approximates it by discretizing the path with a series of lines. Research on reverse ODE solvers aims to precisely estimate the next step, thereby minimizing truncation error from pool discretization in the few-step inference. On the other hand, the discrete method proposed by Karras et al. (2022) does not aim at design a new ODE solving algorithm, but proposes a choice of moving distance $h_{t}$ at each step $t$ to enhance the sample quality. Via the analysis of the Karras’ research, the $h_{t}$ should be large at the beginning of the inference, and then decrease monotonically. Since truncation error is bound by the Equation 5
[…]
where $\mathcal{T}_{i}||\leq Ch_{i}$. The $E$ depends on the discrete step $N$, the start time step $t_{0}$, the ending timestep $t_{N}$, and the existing constant $C$. The meaning of this upper bound is limited by the part that contributes the highest $\mathcal{T}_{i}$. And Karras et al. (2022) reveals that the truncation error of the diffusion model increases monotonously when the $h_{i}$ in each inference step is equal. Based on this observation, Karras et al. (2022) proposes a following discrete method:
Moreover, according to the observation from the Li & Zhang (2025), if we choose $p=7$ and $N=20$, the fourth-to-last step corresponds to the first $\frac{6}{1000}$ perturbation steps, and it will cost $\frac{1}{5}$ inference budget. Based on that observation, Li & Zhang (2025) proposes another discrete method:
When choosing $p=1.2$, the $h_{i}$ in both the noisy part and the less noisy part are larger than those in Karras’s original method. Hence, it can address the problem with the original Karras schedule mentioned above.
@@ 2.3 Diffusion Distillation
The training-free method optimizes the truncation error at each step and thus minimizes the need for additional inference steps for the precise discretization. Unlike this method, the distillation algorithm (Lin et al. (2024); Sauer et al. (2024); Luo et al. ; Song & Dhariwal (2023); Salimans & Ho (2022)) considers the whole inference process of applying a neural network iteratively as a huge neural network. The distillation algorithm, which typically trains a small neural network, reduces its inference steps. However, training a neural network with high-quality content, fast inference, and no additional parameters is difficult. Actually, the distillation algorithm, through the analysis of information theory (Li & Zhang (2025)) and experiments, can sometimes reduce the diversity of the generated images, and affect the FID performance. We do further analysis and experiments to enhance the augmentation.
@@ 2.4 Latent Diffusion Model
Another method for reducing the computational requirements of inference is to encode images into a low-dimensional latent space. Stable diffusion performs inference in a 64 x 64 x 4 latent space to generate a 512 x 512 x 4 image. To achieve this, it firstly encodes the image into a low-dimensional latent space via a $\beta$-variational autoencoder (Burgess et al. (2018)) to let the following inference ignore the perceptual detail. Then, it performs diffusion inference in latent space via a U-Net (Ronneberger et al. (2015)) with an attention mechanism to model the semantic information. Experiments show that compared with vector quantify VAE (Wu & Flierl (2020)), the normal $\beta$-VAE with a small KL-term constraint generates higher-quality images.
@@ 3 Approach
3.1 The effect of the skip-connection modification
Modifying the skip connection via Free-U(Si et al., 2024), a training-free U-Net decorator, can reduce the low-frequency component of the noisy skip connection and enhance the backbone feature, thereby augmenting the denoising capability of the diffusion model while preserving image details(Si et al., 2024). We find that the skip connection modification removes the blur from the generative image, implying a change in the neural network’s original output from a standard score function to a vector pointing to a special image, which allows the model’s behaviour to be similar to that of the diffusion disillation model. However, experiments in Figure 3 show that Free-U will also change the discrete method of the diffusion ODE, and this effect cannot be ignored in the few-step inference that has a poor discrete situation.
Explicitly, the output of the neural network in Equation 3 is pointing to the average of all possible images that can cause this noisy image. In an extreme few-step inference case, like one or two NFE inference, the final output will be blurred, since in that step, the score function only points to the centroid of images. Meanwhile, the distillation algorithm, like rectify flow proposed by Liu (2022), will tend to modify the output and let it point to a special sample.
Although the task of modifying the score function to allow it to point to a special sample without additional training is hard, the Free-U that preserves the detail of the output image while augmenting the denoise capability can partly achieve it in the least noisy stage, since the possible images are very limited. The previous research from Ma et al. (2024) also shows that compared to the original diffusion model, the diffusion distillation model has small skip connection scalers.
However, we don’t know which step could be changed simply from the score function to the conditional score function pointing to special a sample, and one of the vanilla ways is to apply free-U at all time steps, as the original paper claims. This method has been seen as helpful to slightly modify the score function and make it more similar to the distillation’s output. But, via the mathematical analysis of the diffusion ODE and the experimental result in Figure 4. This vanilla method in the extreme few-step inference of high-resolution images, such as 6 NFE, would not work properly.